Circular No. 8517

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 1999gx AND 2005bs Further to IAUC 7258, C. Balland (Université Paris-Sud, Orsay) and

M. Mouchet (Observatoire de Paris, Meudon) report their discovery of another type-Ia supernova in the spectroscopic data set of that 1999 intermediate-z supernova search. SN 1999gx (g′ = 23.3; z = 0.493) was observed on 1999 Sept. 7.08 UT at α = 0^{h}34^{m}15^{s}.47, δ = +0^{o}04′26′′.2 (equinox 2000.0).

Further to IAUC 8516, L. A. G. Monard reports his discovery of an apparent supernova (mag ∼ 17.4 ± 0.3) on unfiltered CCD images taken on Apr. 19.066 UT and confirmed on Apr. 20.014 (at mag ∼ 17.5 ± 0.3). SN 2005bs is located at α = 20^{h}10^{m}14^{s}.11, δ = −56^{o}38′43′′.6 (equinox 2000.0), which is 4′′ west and 15′′ south of the nucleus of the apparent host galaxy (which itself is in the cluster Abell 3667). Nothing is visible at this location on an image from the Digitized Sky Survey (limiting red mag 20.5) or on an image taken on Mar. 24.090 by Monard (limiting red mag 18.5).

SUPERNOVA 2005bo IN NGC 4708 N. Morrell, G. Folatelli, C. Contreras, and M. Hamuy, Las Campanas

Observatory, report that an optical spectrum (range 380–920 nm) of SN 2005bo (cf. IAUC 8514), obtained in the course of the Carnegie Supernova Project on Apr. 18.23 UT with the Las Campanas 2.5-m du Pont telescope (+ WFCCD spectrograph), reveals it to be a normal type-Ia supernova at about maximum light. The expansion velocity derived from the minimum of the Si II 635.5-nm line is 10700 km/s, assuming the recession velocity for NGC 4708 given by Strauss et al. (1992, Ap.J. Suppl. 83, 42).

COMETS C/1997 P4, P5, AND S5 (SOHO) Following are the initial available positions for additional Kreutz sun-

grazing comets found on SOHO website images (continuation to IAUC 8516). K. Battams reports that these three bodies are all “typical C3 Kreutz” objects in appearance — small and stellar.

Comet 1997 UT α_{2000} δ_{2000} Inst. F MPEC
C/1997 P4 Aug. 8.266 8^{h}58.^{m}0 +14^{o}57′ C3 RK 2005-G39
C/1997 P5  8.728 8 59.5 +14 30 C3 RK 2005-G39
C/1997 S5 Sept. 19.636 11 31.7 + 0 36 C3 RK 2005-G39
2005 April 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green